OpenStudy (anonymous):

A nursery operator wants to build a greenhouse in the shape of a half cylinder. The volume of the greenhouse is to be approximately 35,350 cubic feet. A nursery operator wants to build a greenhouse in the shape of a half cylinder. The volume of the greenhouse is to be approximately 35,350 cubic feet. @Mathematics

OpenStudy (anonymous):

do u know the formulae of cylinder

OpenStudy (anonymous):

A. The formula for the radius r (in feet) of a half cylinder is given by \[r=\sqrt{2v/\pi \iota}\]

OpenStudy (anonymous):

where v is the volume (in cubic feet) and l is the length (in feet). Find the radius of the greenhouse. Round the result to the nearest whole number.

OpenStudy (anonymous):

l=100ft

OpenStudy (anonymous):

now plug in all the values in the formulae u will get the answer

OpenStudy (anonymous):

as u know everything so u can do it right !

OpenStudy (anonymous):

I dont understand how to do it

OpenStudy (anonymous):

here, volume = 35350 cu ft l = 100 ft formula for half cylinder radius u had written as \[r = \sqrt(2v/\pi \times l) \]

OpenStudy (anonymous):

\[r = \sqrt{\frac{35350}{3.14 \times 100}}\]

OpenStudy (anonymous):

on solving this see wht u get ??

OpenStudy (anonymous):

I got 15.00530692

OpenStudy (anonymous):

so nearest whole number is 15 ft so radius is 15 ft

OpenStudy (anonymous):

okay, now i have part b & c ; can u help me with this?

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

what is part b say

OpenStudy (anonymous):

wrong formula , one second.

OpenStudy (anonymous):

B. Beams for holding a sprinkler system are to be placed across the top of the greenhouse. The formula for the height h at which the beams are to be placed is given by \[h=\sqrt{r^2-\left(\begin{matrix}a \\ 2\end{matrix}\right)^2}\]

OpenStudy (anonymous):

where a is the length of a beam. rewrite h as a function of only a

OpenStudy (anonymous):

what is a?

OpenStudy (anonymous):

ill upload a picture of the diagram

OpenStudy (anonymous):

\[h = \sqrt{r^2 -(\frac{a}{2})^2}\] square up both the sides what u will get.

OpenStudy (anonymous):

\[h^2 = r^2 -(\frac{a}{2})^2\]

OpenStudy (anonymous):

after that solve for h

OpenStudy (anonymous):

what is given her is r given??

OpenStudy (anonymous):

no r isnt given

OpenStudy (anonymous):

unless r=25

OpenStudy (anonymous):

means unless r = 25???

OpenStudy (anonymous):

no im saying i think r might = 25

OpenStudy (anonymous):

r u have calculated in the last question only so r = 15 ft

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

OMG m also a fool unnecesserily i was showing those things just plug in the value of r in the equation which u gave me for h

OpenStudy (anonymous):

lol its okay